what is the exact value of the solution to this equation? 32.4=(68.4)r^12
Found 2 solutions by shree840, stanbon:
Answer by shree840(260)

32.4=68.4r^12
or r^12=32.4/68.4=.473682
or r=.939654
( I used the calculator from the accessories)

Answer by stanbon(75887)

what is the exact value of the solution to this equation?
32.4=(68.4)r^12
---
Divide to get:
r^12 = 32.4/68.4 = 0.473684
Take the 12th root to get:
r = 0.939531...
Cheers,
Stan H.
=========